### v2

# landmarker: a Toolkit for Anatomical Landmark Localization in 2D/3D Images

Abstract
Anatomical landmark localization in 2D/3D images is a critical task in medical imaging. Although many general-purpose tools exist for landmark localization in classical computer vision tasks, such as pose estimation, they lack the specialized features and modularity necessary for anatomical landmark localization applications in the medical domain. Therefore, we introduce landmarker, a Python package built on PyTorch. The package provides a comprehensive, flexible toolkit for developing and evaluating landmark localization algorithms, supporting a range of methodologies, including static and adaptive heatmap regression. landmarker enhances the accuracy of landmark identification, streamlines research and development processes, and supports various image formats and preprocessing pipelines. Its modular design allows users to customize and extend the toolkit for specific datasets and applications, accelerating innovation in medical imaging. landmarker addresses a critical need for precision and customization in landmark localization tasks not adequately met by existing general-purpose pose estimation tools.
Code: https://github.com/predict-idlab/landmarker.
Published [^1] Published in SoftwareX: https://doi.org/10.1016/j.softx.2025.102165
Keywords Landmark localization  $\cdot$ Keypoint detection  $\cdot$ Medical image analysis  $\cdot$ Python  $\cdot$ PyTorch
1 Motivation and significance
Landmark (or keypoint) localization in 2D/3D images is a fundamental challenge in computer vision, crucial for tasks such as pose estimation Luvizon et al. (2018); Tompson et al. (2014); Newell et al. (2016), face alignment Zhou et al. (2023); Kumar et al. (2019), robotic manipulation Ziegler et al. (2020), surgical procedures Sharan et al. (2021), image registration (Grewal et al., 2023), various image-based medical diagnoses de Queiroz Tavares Borges Mesquita et al. (2023); Ryu et al. (2022); Pei et al. (2023); Payer et al. (2019); Thaler et al. (2021) and treatment response modeling in longitudinal images (Ou et al., 2015; Cai et al., 2021). The effectiveness of different approaches is often domain and task-dependent, necessitating extensive experimentation for new applications. Popular methods, such as heatmap regression, involve numerous preprocessing and post-processing steps. The latter, such as sub-pixel accuracy methods, can significantly improve the precision needed in medical applications but nevertheless are frequently overlooked.
As existing packages such as OpenPose Cao et al. (2021) and Ultralytics Jocher et al. (2023) are primarily designed for pose estimation, they lack the modularity needed for medical applications. While MMPose MMPose Contributors (2020) offers different components for various tasks, it is still tailored toward pose estimation and does not easily integrate with other frameworks or handle specific medical imaging requirements. Although MMPose (MMPose Contributors, 2020) supports 3D landmark localization, it cannot process 3D input data, a critical requirement for medical imaging applications. Furthermore, while MMPose offers some flexibility in model creation, its pipeline mandates the use of its framework for preprocessing, training, and evaluation, limiting compatibility with deep learning frameworks such as PyTorch Lightning (Falcon and The PyTorch Lightning team, 2019). Additionally, neither pose estimation library supports medical imaging formats (e.g., NIfTI and DICOM) or provides specialized data augmentation techniques to simulate common acquisition artifacts. Table 1 summarizes the key differences between these existing packages and the proposed landmark detection package within the medical imaging context.
3D Medical
Package Custom Keypoints Images Modular Format Aug.
ultralytics (Jocher et al., 2023)
MMPose (MMPose Contributors, 2020)
landmarker
Table 1: Comparison of Python packages for keypoint/landmark localization.
Our Python package, landmarker, addresses the need for precise anatomical landmark localization in medical images, which is critical for diagnostics and therapeutic procedures in different specialties, such as orthodontics, maxillofacial surgery, and orthopedics. It is a flexible PyTorch-based toolkit designed explicitly for training, modeling, and evaluating anatomical landmark localization problems. This toolkit accelerates the development of algorithms and enhances the accuracy of landmark identification in medical images, thereby improving diagnostic and treatment outcomes. Its modular and adaptable framework allows researchers and practitioners to implement state-of-the-art algorithms tailored to their specific datasets.
landmarker has a modular framework, which supports various landmark localization algorithms. This flexibility enables customization and extension according to specific needs. The package offers interfaces for data preprocessing, model training, evaluation, and visualization.
2 Background: landmark localization
There are two primary methodologies in landmark localization (see Figure 1): heatmap regression and coordinate regression. Coordinate regression (Fard and Mahoor, 2022) involves learning a direct relationship between the image and the landmark locations. However, the introduction of heatmap regression Tompson et al. (2014), which generally yields better performance, has shifted the research focus. Heatmap regression
Figure 1: Taxonomy of (deep learning) landmark localization approaches. The frequently used taxonomy in the literature is highlighted in yellow, while our extended taxonomy of the problem is highlighted in blue.
predicts a heatmap that assigns a likelihood or probability to each (downsized) image pixel. Early heatmap regression methods, termed direct heatmap regression, assume a parametric distribution, typically a bivariate Gaussian or Laplacian distribution. Direct heatmap regression can be further classified into static and adaptive approaches. In static heatmap regression, the heatmap distribution parameters are fixed hyperparameters during training. In adaptive heatmap regression, these parameters are adjusted during training, either by treating them as learnable model parameters (Payer et al., 2019, 2020; Thaler et al., 2021) or by using scheduling methods that modify the parameters based on specific evaluation metrics (Teixeira et al., 2019). Another variant of static heatmap regression involves training a model on one-hot heatmap images (masks), essentially parameterizing the heatmap as a categorical distribution and transforming the task into a multi-class classification problem where each pixel represents a class (McCouat and Voiculescu, 2022).
Recent advancements on heatmap regression have seen the rise of fully convolutional neural networks (CNNs) with differentiable decoding operations, such as the soft-argmax operation Luvizon et al. (2018). These methods, prominent in facial landmark localization, use the heatmap generation layer as an intermediate layer and the decoding operation as the final layer, optimizing the network with a loss function that compares the decoded coordinates to the ground-truth coordinates. Some of these methods account for the ambiguity in landmark ground truth by incorporating uncertainty-aware loss functions, such as the Gaussian log-likelihood loss (Kumar et al., 2019), or other custom loss functions Kumar et al. (2019, 2020); Zhou et al. (2023).
In the medical field, landmark localization often utilizes two-stage approaches (Jiang et al., 2022; Song et al., 2020; Zhong et al., 2019; Nguyen et al., 2020). The first stage identifies landmarks on a low-resolution image, while the second stage refines this localization using a high-resolution patch, which is a region of interest based on the initial predictions. One of the previously described approaches can be applied in both stages. Another technique, also used in medical imaging, involves segmenting or contouring specific anatomical structures and inferring the landmarks from the resulting masks or contours (Gooßen et al., 2011; Pei et al., 2021).
The diversity of approaches in landmark localization, from basic coordinate regression to sophisticated uncertainty-aware heatmap methods, highlights the complexity of this field, particularly in medical imaging applications. While general-purpose tools like OpenPose and MMPose exist, they primarily focus on human pose estimation and lack the specialized features needed for medical imaging applications. These tools often do not support the full range of heatmap generation and decoding methods, uncertainty quantification, or medical image-specific preprocessing pipelines crucial for anatomical landmark localization. Additionally, existing tools typically do not provide the modularity required to implement and experiment with different approaches, such as two-stage refinement methods commonly used in medical applications. These limitations, combined with the need for high precision in medical contexts, underscore the need for a specialized toolkit to handle the unique challenges of anatomical landmark localization while supporting established and emerging methodologies.
3 Software description
landmarker is a Python package that leverages PyTorch Paszke et al. (2019) deep learning framework and MONAI Cardoso et al. (2022), a PyTorch-based deep learning framework for healthcare imaging for handling medical image files and transformations.
Users can install landmarker via pip. After installation, our documentation and variety of examples will guide users in applying the toolkit for their specific needs.
3.1 Software architecture and functionalities
Figure 2: Flowchart of the different sub-packages of the landmarker pacakge.
The landmarker package is structured into several key modules, each addressing specific aspects of the landmark localization pipeline, see Figure 2.
3.1.1 Data handling and preprocessing
The data module provides flexible "dataset" classes that inherit from PyTorch’s Dataset, ensuring compatibility with PyTorch DataLoaders. Four main dataset types are supported:
  • LandmarkDataset: For images with corresponding landmark coordinates.
  • HeatmapDataset: For images with corresponding static heatmaps representing landmarks.
  • MaskDataset: For images with binary segmentation masks indicating landmark locations.
  • PatchDataset: For image patches or regions of interest (ROI) with corresponding landmarks.
These datasets support various image formats (NIfTI, DICOM, PNG, JPG, BMP, NPY/NPZ) through integration with the MONAI library (Cardoso et al., 2022). The module can handle single-class, multi-class single-instance, and multi-class multi-instance landmark problems. Utility functions are provided to transform common annotation formats (e.g., labelme Wada (2024)) into the required format.
The package also includes the datasets module, which contains functions to import benchmark datasets, such as the ISBI2015 Wang et al. (2016) dataset directly. Preprocessing and data augmentation capabilities are available, leveraging MONAI’s transformations for 2D and 3D data. The package handles affine landmarks transformations through the transforms module.
3.1.2 Heatmap generation and decoding
The heatmap module provides functionality for generating target heatmaps and decoding predicted heatmaps, which is essential for heatmap regression approaches.
3.1.2.1 Heatmap generation
The HeatmapGenerator class and its subclasses support various parametric distributions (e.g., multivariate Gaussian and Laplace). The design allows easy implementation of custom distributions. The parameters of the parametric distribution can be set on initialization. Still, they can also be changed during training, for example, for adaptive heatmap regression approaches that rely on a scheduler, such as Adaloss scheduler (Teixeira et al., 2019). Additionally, the parameters could also be learnable parameters during a training procedure.
3.1.2.2 Heatmap decoding
While heatmap generation is only needed for direct approaches, decoding heatmaps is needed for all heatmap regression approaches, with the caveat that the indirect approaches need differentiable decoding operations. Multiple decoding methods are implemented:
The argmax operation is the simplest decoder operation Tompson et al. (2014). The operation takes the maximum value of the heatmap $H_{i}$,
$\hat{y}_{i}=\text{argmax}H_{i},$ (1)
to get the coordinate $\hat{y}_{i}$ of landmark $i$. The major downside of this approach is that it introduces a discretization error by only choosing a pixel as the predicted landmark.
The weighted spatial mean (Luvizon et al., 2018; Dong et al., 2018; Kumar et al., 2019, 2020) is the approach for indirect heatmap generation. The operation calculates the landmark location $\hat{y}_{i}$ by first post-processing the heatmap $H_{i}$ with an activation function $\sigma$, typically a type of normalization operation, and afterward taking the weighted spatial mean, e.g., the two-dimensional case:
$\hat{y}_{i}=\left(\sum_{c=0}^{W}\sum_{l=0}^{H}\frac{c}{W}\sigma(H_{i})_{l,c},\sum_{c=0}^{W}\sum_{l=0}^{H}\frac{l}{H}\sigma(H_{i})_{l,c}\right)$ (2)
where $W$ and $H$ are, respectively, the heatmap’s width and height. Frequently used activation functions $\sigma$ are a ReLU activation combined with a normalization procedure and the softmax operation on the heatmap, which is in the literature often referred to as soft-argmax operation (Luvizon et al., 2018; Dong et al., 2018; Bulat et al., 2021).
A significant issue of taking the weighted spatial mean is that it can globally lead to semantically unstructured outputs and thus decrease performance (Bulat et al., 2021). In Bulat et al. (2021), they propose to apply soft-argmax operation locally, i.e., a small window, around the heatmap location with a maximum heatmap value. We implement this method and extend it to be used with any post-processing function, i.e., local weight spatial mean operation.
3.1.3 Models and losses
While landmarker is compatible with any PyTorch model or loss function, it includes implementations of successful approaches from the literature, such as the spatial configuration network Payer et al. (2019), which is effective for positionally consistent medical images like cephalograms. Also, multiple heatmap regression loss functions are implemented from successful approaches (Thaler et al., 2021; Nibali et al., 2018; Kumar et al., 2019; Feng et al., 2018; Wang et al., 2019; Zhou et al., 2023).
3.1.4 Evaluation and visualization
The visualize module allows inspecting constructed datasets and trained models. Additionally, it enables the generation of detection reports that output several metrics from the metrics module, such as point error and success detection rate (SDR), which is the percentage of predicted landmarks with a point error smaller or equal to a specified radius.
4 Illustrative examples
In this section, we illustrate the use of landmarker. We will perform adaptive heatmap regression with the ISBI2015 cephalometric landmark dataset, showcasing most functionality.
We start by loading the data into a LandmarkDataset by providing a list of paths to images, a NumPy array of the shape $(N,C,D)$ where $N$ is the number of samples, $C$ is the number of landmark classes, and $D$ represents the spatial dimension which can be 2 or 3.
1from landmarker.data import LandmarkDataset
2
3image_paths = ... # list of paths of images
4landmarks = ... # numpy array or torch tensor provided
5pixel_spacing = ... # pixel spacing of the images
6class_names = ... # names of the landmarks
7compose_transform = ... # MONAI compose transform
8
9ds = LandmarkDataset(
10 imgs=image_paths,
11 landmarks=landmarks,
12 spatial_dims=2,
13 transform=compose_transform,
14 dim_img=(512,512) # dimension to resize to
15 class_names=class_names)
Listing 1: Loading data into a LandmarkDataset
Alternatively, one can also use the dataset module, which includes everything to import the ISBI2015 dataset directly.
1from landmarker.dataset import get_cepha_landmark_datasets
2
3data_dir = ...
4ds_train, ds_test1, ds_test2 = get_cepha_landmark_datasets(data_dir)
Listing 2: Loading data into a LandmarkDataset through the dataset module.
Once the data is loaded in the proper format, we can set up the heatmap generator to generate a heatmap from landmarks.
1from landmarker.heatmap import GaussianHeatmapGenerator
2
3generator = GaussianHeatmapGenerator(
4 nb_landmarks=19, # number of landmarks
5 sigmas=3, # sd. value of the Gaussian heatmap function
6 learnable=True, # make the covariance matrix learnable parameters
7 heatmap_size=(512,512))
Listing 3: Setting up Gaussian heatmap generator.
After this, the user can visually inspect if everything is initialized as expected by using the inspection_plot function (Listing 4 and Figure 3).
1from landmarker.visualize import inspection_plot
2
3inspection_plot(ds_train, range(3), heatmap_generator=heatmap_generator)
Listing 4: Loading data into a LandmarkDataset
Figure 3: Example of results of running the inspection_plot on the ISBI2015 dataset.
With the LandmarkDataset and HeatmapGenerator adequately set up, we can now train a model to output likelihood heatmaps. In Listing 5, we train a SpatialConfigurationNetwork Payer et al. (2019); Thaler et al. (2021) and parameterization of the Gaussian heatmap Payer et al. (2020); Thaler et al. (2021).
1import torch
2
3from landmarker.models import OriginalSpatialConfigurationNet
4from landmarker.losses import GaussianHeatmapL2Loss
5
6model = OriginalSpatialConfigurationNet(in_channels=1, out_channels=19)
7optimizer = torch.optim.SGD([
8{’params’: model.parameters(), "weight_decay":1e-3},
9{’params’: heatmap_generator.sigmas},
10{’params’: heatmap_generator.rotation}]
11, lr=1e-6, momentum=0.99, nesterov=True)
12
13criterion = GaussianHeatmapL2Loss(alpha=5)
14
15train_loader = DataLoader(ds_train, batch_size=1, shuffle=True, num_workers=0)
16
17# Start training
18for epoch in range(100):
19 model.train()
20 for batch in train_loader:
21 images = batch["image"].to(device)
22 landmarks = batch["landmark"].to(device)
23 optimizer.zero_grad()
24 outputs = model(images)
25 heatmaps = heatmap_generator(landmarks)
26 loss = criterion(outputs, heatmap_generator.sigmas, heatmaps)
27 loss.backward()
28 optimizer.step()
Listing 5: Training adaptive heatmap regression model.
After the training procedure, one can visually evaluate the results by using the prediction_inspect_plot function (see Listing 6 and Figure 4) or by creating a detection report.
1from landmarker.visualize import prediction_inspect_plot
2
3prediction_inspect_plot(ds_test1, model, ds_test1.indices[:3])
Listing 6: Prediction inspection plot.
(a) ISBI2015 dataset
(b) Pelvis X-rays
(c) Endoscopic images
Figure 4: Examples of landmark predictions using the prediction_inspection_plot tool for different datasets: (a) Endoscopic images, (b) ISBI2015, and (c) Pelvis X-rays.
4.1 Benchmarks
To evaluate our package capabilities and its model’s landmark localization precision quantitatively, we conducted two benchmark experiments on relatively recent datasets that showcase its performance on both 2D and 3D medical imaging tasks. For the 2D benchmark, we used pelvis x-ray images from the Osteoarthritis Initiative (Pei et al., 2023). In this experiment, our spatial configuration network (SCN) with adaptive heatmaps—implemented in landmarker, was compared against a state-of-the-art U-Net with Attention approach (Pei et al., 2023). As shown in Table 2, our method achieved a lower average position error (PE) and higher success detection rates (SDR) across multiple thresholds, clearly demonstrating the capabilities of the package as a benchmarking tool.
Table 2: Landmark localization results of spatial configuration network (SCN) with adaptive heatmaps (Payer et al., 2020; Thaler et al., 2021) (implemented in landmarker) on the pelvis x-rays (2D) of the osteoarthritis initiative dataset (Pei et al., 2023) compared against the literature. The best results are highlighted in bold. A 5-fold validation approach, as suggested by Pei et al. (Pei et al., 2023), obtains the results.
Model name PE (mm) SDR (%)
1 mm 2 mm 3 mm 4 mm
U-Net w/ Attention (Pei et al., 2023)
3.14 8.65% 33.89% 59.25% 78.61%
SCN (Payer et al., 2020; Thaler et al., 2021)
1.61 33.25% 74.38% 90.96% 95.85%
In addition to the 2D evaluation, we benchmarked our model on 3D images using CT scans of the skull (He et al., 2024), where the task was to annotate mandibular landmarks automatically. The results, presented in Table 3, include both validation and test metrics, with our one-hot ensemble (Jonkers et al., 2025) method outperforming the competitive baseline regarding PE and SDR. These results confirm our model’s precision and highlight its flexibility in handling diverse anatomical landmarking applications across different imaging modalities.
Table 3: Landmark localization results of one hot ensemble (Jonkers et al., 2025) (implemented in landmarker) on the MML (3D) dataset compared against the literature. This benchmark dataset is the same as the data subset, only with complete landmarks, used as a benchmark in He et al. (He et al., 2024). The best results are highlighted in bold.
Model name Validation Test
PE (mm) SDR (%) PE (mm) SDR (%)
2 mm 2.5 mm 3 mm 4 mm 2 mm 2.5 mm 3 mm 4 mm
Pruning-ResUNet3D (He et al., 2024)
1.82 73.21% 82.14% 88.93% 94.76% 1.96 70.03% 79.97% 86.10% 92.73%
One hot ensemble (Jonkers et al., 2025)
1.60 77.81% 87.76% 89.92% 93.37% 1.39 81.67% 91.31% 93.33% 96.31%
5 Impact and conclusions
The development and release of the landmarker Python package offer significant advancements in anatomical landmark localization, particularly within medical imaging. By providing a flexible and modular toolkit tailored explicitly for medical imaging applications, landmarker addresses a critical need for precision and customization in landmark localization tasks that are not adequately met by existing general-purpose pose estimation tools like OpenPose or MMPose.
One of the key impacts of landmarker is its ability to streamline the research and development process for researchers and developers working on image-based diagnostic and therapeutic deep learning applications. The package’s integration with PyTorch and MONAI ensures compatibility with various medical image formats and processing pipelines, facilitating easy adoption in research environments. Moreover, the extensive support for different landmark localization methodologies, including static and adaptive heatmap regression, allows users to implement state-of-the-art algorithms with minimal overhead.
Including customizable modules for data handling, preprocessing, heatmap generation, and model evaluation further enhances the utility of landmarker. These features enable users to quickly adapt the package to their specific use cases, whether they involve 2D or 3D images, single-class or multi-class landmark localization, or static versus adaptive methodologies. The ability to implement and test novel approaches within a unified framework accelerates innovation in computer-aided diagnosis through medical imaging, potentially leading to more accurate diagnoses and better patient outcomes.
In conclusion, landmarker represents a powerful tool for advancing the field of anatomical landmark localization in medical imaging. Its flexibility and ease of use make it a helpful resource for researchers and engineers. We plan to incorporate future developments and improvements, particularly in uncertainty quantification, which is critical for enhancing the reliability and robustness of landmark localization in medical applications.
CRedit Authorship Contribution Statement
Jef Jonkers: Conceptualization, Methodology, Software, Validation, Visualization, Writing - Original Draft. Luc Duchateau: Conceptualization, Writing - Review & Editing, Supervision, Funding acquisition. Glenn Van Wallendael: Conceptualization, Writing - Review & Editing, Supervision. Sofie Van Hoecke Conceptualization, Writing - Review & Editing, Supervision, Funding acquisition.
Funding sources
Jef Jonkers is funded by the Research Foundation Flanders (FWO, Ref. 1S11525N). This research was (partially) funded by the Flemish Government (AI Research Program).